OpenStudy (anonymous):

can anyone help me with this problem.. 6x + 2y = –88

OpenStudy (anonymous):

what do you want to do with it ?

OpenStudy (anonymous):

y=-3x-44 thats as far as i could take it

OpenStudy (anonymous):

What answer are you looking for?

OpenStudy (anonymous):

Find the slope and y-intercept of the line

OpenStudy (anonymous):

take what @babygirl11271963 did which is in the form of y = mx+b m is the slope b is the y-intercept

OpenStudy (anonymous):

but how would I do that im new to this problem...????

OpenStudy (anonymous):

@Philippians413

OpenStudy (anonymous):

slope is -3 and y intercept is -44

OpenStudy (anonymous):

put it in slope intercept form of y = mx+b m being the slope b being the y-intercept you have \[6x+2y=-88 \] subtract 6x to both sides \[2y= -6x-88 \] divide everything by to to isolate y \[y = -3x-44 \]

OpenStudy (anonymous):

i took 6x+2y=-88 and subtracted 6x from both sides to move it to the opposite side. now you have 2y=-6x-88. divide it all by 2 so you can single out the y. you get y=-3x-44. y and x are your variables and your slope is -3 and you y intercept is -44

OpenStudy (anonymous):

yeah what he said XD

OpenStudy (anonymous):

okay so this how you do it 6x+2y=-88 get 6x on the other side of the equation so 2y=-6x-88 then for the intercept form u need to get rid of the 2 in front of the y so the equation would end in y=-6/2-88/2 so y=-3-44 You see?

OpenStudy (anonymous):

thank you to al of you :) this was a a lot of help with my future problems.. :) wish I can give all of you guys medals

OpenStudy (anonymous):

now you have it in slope intercept form \[y = mx+b \] \[y = \color {red}{-3}x\color{red}{-44}\]

OpenStudy (anonymous):

we all saif the saem thing :| i think he gets it

OpenStudy (anonymous):

said same ****

OpenStudy (anonymous):

lol yea I do

OpenStudy (anonymous):

you are welcome glad to help!!

OpenStudy (anonymous):

welcome hun :)

OpenStudy (anonymous):

can someone help me with this Find the slope of the line that passes through the pair of points.

OpenStudy (anonymous):

(2, 6), (7, 0)